OpenStudy (anonymous):

In a sentence, compare the amplitudes of the two functions: F(x)= 2cos(pix) G(x)= 1/2cos(2x) II. In a sentence, compare the periods of the two functions. Which function has the larger period?

OpenStudy (anonymous):

the first one has amplitude 2, the second has amplitude \(\frac{1}{2}\) the amplitude is the absolute value of the coefficient out front

OpenStudy (anonymous):

the period of \(\sin(bx)\) is \(\frac{2\pi}{b}\) so the period of the first one is \(\frac{2\pi}{\pi}=2\) and the period of the second one is \(\frac{2\pi}{2}=\pi\) you can figure out what is larger, \(2\) or \(\pi\)

OpenStudy (anonymous):

ok it says "cosine" but the same reasoning applies